The area of a triangle is 105 centimeters. If the base is 14 centimeter. What is the height of the triangle
Answer by ptaylor(1894)

The area of a triangle is 105 centimeters. If the base is 14 centimeter. What is the height of the triangle
Area of a Triangle = (1/2)Base(B) times Height(H) or A=(1/2)BH
A=105 sq cm
B=14 cm

105=(1/2)(14)H
105=7H
H=15 cm-------------------------answer

Hope this helps----ptaylor